# Structurally damped semilinear evolution equation for positive operators on Hilbert space

### v1

@@ 5. Proof of main results
$\partial_{t}^{k}\widehat{E}_{1}(t,\xi)=\frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)^{k}}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)t}-\frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)^{k}}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)t},$ (5.25)
@@
where we used the argument similar to the estimate of $\partial_{t}^{k}\widehat{E}_{1}$ on $\mathcal{R}_{2}$ and consequently, using the relation (4.9), we obtain
@@
In contrast to the previous instance, where exponential decay was accomplished at the expense of having Sobolev regularity on the initial data, in this situation we are able to eliminate the $\lambda_{\xi}$ factor in the region $\mathcal{R}_{4}$ by using the relation (4.12). The decay rate is weaker with this strategy-exhibiting singular behavior near $t=0$. However, we obtain this weak decay estimates for initial data with bare minimal regularity i.e., $(u_{0},u_{1})\in\mathcal{H}$ only. Similar behavior will also be observed in the the case other damping scenarios.
Having established the decay estimates in the case of effective damping, we can prove Theorem 2.6 for the case $2\theta\in(0,\sigma)$, with some modifications described in the following remark.
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda^{\beta}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}\lesssim e^{-\delta t}\quad\text{and}\quad|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim e^{-\delta t}.$
@@
$|\widehat{E}_{1}(t,\xi)|\lesssim\frac{1}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{1}{2}\sqrt{4\lambda_{\xi}^{\sigma}-\lambda_{\xi}^{2\theta}}\right)t}\lesssim e^{-\lambda_{\xi}^{\sigma-\theta}t}\lesssim e^{-\delta t}.$
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\lambda_{\xi}^{\sigma-\theta}t}=\left(\lambda_{\xi}^{\sigma-\theta}t\right)^{\frac{\beta}{\sigma-\theta}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}t^{-\frac{\beta}{\sigma-\theta}}\lesssim\begin{cases}t^{-\frac{\beta}{\sigma-\theta}}&\text{if }t\in(0,1],\\ e^{-\delta t}&\text{if }t\in[1,\infty),\end{cases}$ (5.61)
@@
and hence the estimate (5.68) and (5.69) will be replaced by (2.5). This completes the proof of Theorem 2.6 for the case $2\theta\in(\sigma,2\sigma]$.
@@
The proof of Theorem 2.6 is now complete, as demonstrated in Remarks 5.3, 5.4, and 5.5. The estimates in the high-frequency region are alternatively derived in these remarks to obtain exponential decay in instead of polynomial decay. Nevertheless, there are certain terms $\mathcal{R}_{1}$ in $\|u\|_{\mathcal{H}}$ and $\|\partial_{t}u\|_{\mathcal{H}}$ that do not show any decay.
@@ 6. Application
Having established the decay estimates for both the undamped and damped cases of the semilinear Cauchy problem (1.3), we are now able to employ these results to the question of global (in-time) well-posedness. This application emphasizes the decay framework’s efficiency in managing nonlinear effects and guaranteeing the long-term existence and uniqueness of solutions under appropriate assumptions regarding the initial data and the nonlinearity.
@@
Let $\sigma>0$, $\theta\in[0,\sigma]\,$ and $\mathcal{L}$ be strictly positive operator. Let $p$ and $f$ satisfies the following relation:
$\|f(u)-f(v)\|_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}\lesssim(\|u\|^{p-1}_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}+\|v\|^{p-1}_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}})\|u-v\|_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}\quad\text{and }\quad f(0)=0.$
Also assume that $(u_{0},u_{1})\in\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}\times\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}$ where
  1. (1)
    $m_{k,\beta}=k\sigma+2\beta$ and $n_{k,\beta}=\max\{(k-1)\sigma,0\}+\begin{cases}2\beta-\sigma\quad&\text{if }2\beta>\sigma,\\ 0&\text{if }2\beta\leq\sigma,\end{cases}$ for $2\theta\in[0,\sigma]$ and
  2. (2)
    $m_{k,\beta}=2k\theta+2\beta$ and $n_{k,\beta}=\max\{2(k-1)\theta,0\}+\begin{cases}2\beta-\sigma\quad&\text{if }2\beta>\sigma,\\ 0&\text{if }2\beta\sigma,\end{cases}$ for $2\theta\in(\sigma,2\sigma]$.
@@
$(u_{0},u_{1})\in\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}\times\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}~{}~{}\text{ satisfying }~{}~{}\|u_{0}\|_{\mathcal{H}^{m_{k,\beta}}_{\mathcal{L}}}+\|u_{1}\|_{\mathcal{H}^{n_{k,\beta}}_{\mathcal{L}}}<\epsilon,$
there is a unique global solution $u\in C^{k}(\mathbb{R}_{+};\mathcal{H}_{\mathcal{L}}^{\beta})$ to the Cauchy problem (1.3). Moreover, u satisfies the following estimates
$\left\|\partial_{t}^{i}\mathcal{L}^{j}u(t,\cdot)\right\|_{\mathcal{H}}\lesssim e^{-\delta t}\left(\|u_{0}\|_{\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}}+\|u_{0}\|_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}\right),$
for $(i,j)\in\{(0,0),(0,\beta),(1,0)\}$ and for all $t>0$.
@@
represents the solution of the homogeneous Cauchy problem (2.1), where $*_{\mathcal{L}}$ denotes $\mathcal{L}$-convolution. The solution to the semilinear Cauchy problem (1.3), using Duhamel’s principle, can be expressed as
@@
Now, we introduce the solution space $X^{k,\beta}:=C^{k}(\mathbb{R}_{+};\mathcal{H}^{2\beta})$ associated with the norm
$\left\|u\right\|_{X^{k,\beta}}:=\sup\limits_{t\in[0,\infty)}\left\{e^{\delta t}\left(\left\|u(t)\right\|_{\mathcal{H}}+\left\|\mathcal{L}^{\beta}u(t)\right\|_{\mathcal{H}}+\left\|\partial^{k}_{t}u(t)\right\|_{\mathcal{H}}\right)\right\}.$
Let $\mathcal{N}:X^{k,\beta}\to X^{k,\beta}$ be the operator defined as follows:

### v2

@@ 5. Proof of main results
$\partial_{t}^{k}\widehat{E}_{1}(t,\xi)=\frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)^{k}}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)t}-\frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)^{k}}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}\right)t}$
@@
where we used the argument similar to the estimate of $\partial_{t}^{k}\widehat{E}_{1}$ on $\mathcal{R}_{2}$ in Theorem 2.1 and consequently, using the relation (4.10), we obtain
@@
In contrast to the previous instance, where exponential decay was accomplished with Sobolev regularity on the initial data, in this situation we are able to eliminate the $\lambda_{\xi}$ factor in the region $\mathcal{R}_{4}$ by using the relation (3.4).
Having established the decay estimates in the case of effective damping, we can now obtain improved decay estimates with Sobolev regularity on the initial Cauchy data, as stated in Theorem 2.4 for effective damping, with some modifications outlined in the following remark.
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}\lesssim e^{-\delta t}\quad\text{and}\quad|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim e^{-\delta t}.$
@@
$|\widehat{E}_{1}(t,\xi)|\lesssim\frac{1}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\sigma}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\sigma}}\right)t}\lesssim e^{-\lambda_{\xi}^{\sigma-\theta}t}\lesssim e^{-\delta t}.$
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\lambda_{\xi}^{\sigma-\theta}t}=\left(\lambda_{\xi}^{\sigma-\theta}t\right)^{\frac{\beta}{\sigma-\theta}}e^{-\lambda_{\xi}^{\sigma-\theta}t}t^{-\frac{\beta}{\sigma-\theta}}\lesssim\begin{cases}t^{-\frac{\beta}{\sigma-\theta}}&\text{if }t\in(0,1],\\ e^{-\delta t}&\text{if }t\in[1,\infty),\end{cases}$ (5.60)
@@
and hence the estimate (5.67) and (5.68) will be replaced by (2.5). This completes the proof of Theorem 2.4 for the case $2\theta\in(\sigma,2\sigma)$.
@@
The proof of Theorem 2.4 for effective damping, critical damping and non-effective damping is now complete, as demonstrated in Remarks 5.3, 5.4, and 5.5. The estimates in the high-frequency region are alternatively derived in these remarks to obtain exponential decay in instead of polynomial decay. Nevertheless, there are certain terms $\mathcal{R}_{1}$ in $\|u\|_{\mathcal{H}}$ and $\|\partial_{t}u\|_{\mathcal{H}}$ that do not show any decay.
Let us now proceed to prove the decay estimates for viscoelastic type damping in Theorem 2.4.
Proof.
Similar to the previous cases, it is reasonable to consider the following partition of $\mathcal{I}$ in this case:
  1. (1)
    $\mathcal{R}_{1}=\{\xi\in\mathcal{I}:\lambda_{\xi}^{\theta}=0\},$
  2. (2)
    $\mathcal{R}_{2}=\left\{\xi\in\mathcal{I}:0<\lambda_{\xi}^{\theta}<4\right\}$,
  3. (3)
    $\mathcal{R}_{3}=\left\{\xi\in\mathcal{I}:\lambda_{\xi}^{\theta}=4\right\}$, and
  4. (4)
    $\mathcal{R}_{4}=\left\{\xi\in\mathcal{I}:\lambda_{\xi}^{\theta}>4\right\}$.
Combining the relation (4.6) with representation (4.8), we get
$\widehat{E}_{0}(t,\xi)=\begin{cases}1&\text{if }\xi\in\mathcal{R}_{1},\\ \left[\cos\left(\frac{1}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}t\right)+\frac{\lambda_{\xi}^{\theta}\sin\left(\frac{1}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}t\right)}{\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ \left[1+\frac{1}{2}\lambda_{\xi}^{\theta}t\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \left[\cosh\left(\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}t\right)+\frac{\lambda_{\xi}^{\theta}\sinh\left(\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}t\right)}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4},\end{cases}$ (5.69)
and
$\widehat{E}_{1}(t,\xi)=\begin{cases}t&\text{if }\xi\in\mathcal{R}_{1},\\ \frac{\sin\left(\frac{1}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}t\right)}{\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ te^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \frac{\sinh\left(\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}t\right)}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4}.\end{cases}$ (5.70)
Estimate on $\mathcal{R}_{2}$: From (5.69) and (5.70), using the fact that $\frac{\lambda_{\xi}^{\theta}}{\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}$ is bounded and $\left|\sin x\right|\leq 1$, we get
Further, for any $\beta>0$ we have
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}\lesssim e^{-\delta t}\quad\text{and}\quad|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim e^{-\delta t}.$
Consequently, we can estimate
$\partial_{t}^{k}\widehat{E}_{1}(t,\xi)=\\ \frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)^{k}}{i\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)t}-\frac{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)^{k}}{i\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)t}.$ (5.73)
Using the boundedness of $\frac{1}{\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}}$, we obtain
$|\partial_{t}^{k}\widehat{E}_{1}(t,\xi)|$ $\lesssim\left(\left|\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)^{k}\right|+\left|\left(-\frac{1}{2}\lambda_{\xi}^{\theta}-\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)^{k}\right|\right)e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}\lesssim e^{-\delta t},$
Consequently using (4.10), we get
$|\partial_{t}^{k}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\theta}|\partial_{t}^{k-1}\widehat{E}_{1}(t,\xi)|\lesssim\lambda_{\xi}^{\frac{k+1}{2}\theta}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}\lesssim e^{-\delta t}.$
Hence we obtain
Estimate on $\mathcal{R}_{3}$: From (5.69) and (5.70), we have
This gives
Consequently, for any $\beta>0$,
Furthermore, using the relation (4.11) and (5.70), we have
and consequently, using the relation (4.10), we obtain
Combining the above estimate with the representation (4.9), we have
Estimate on $\mathcal{R}_{4}$: From (5.69) and (5.70), we have
$|\widehat{E}_{0}(t,\xi)|\lesssim\left[1+\frac{\lambda_{\xi}^{\theta}}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}}\right]e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}\right)t}\lesssim e^{-\lambda_{\xi}^{\theta-\theta}t}\lesssim e^{-t},$
and
$|\widehat{E}_{1}(t,\xi)|\lesssim\frac{1}{\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}}e^{\left(-\frac{1}{2}\lambda_{\xi}^{\theta}+\frac{1}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}\right)t}\lesssim e^{-\lambda_{\xi}^{\theta-\theta}t}\lesssim e^{-t}.$
$|\widehat{u}(t,\xi)|\lesssim e^{-t}\left(|\widehat{u}_{0}(\xi)|+|\widehat{u}_{1}(\xi)|\right).$ (5.78)
Further, for any $\beta>0$,
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-t}$ (5.79)
and
$|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim\lambda_{\xi}^{\beta-\frac{\theta}{2}}e^{-t},$ (5.80)
whenever $2\beta>\theta$, while
$|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim\lambda_{\xi}^{\beta-\frac{\theta}{2}}e^{-t}\lesssim e^{-t},$ (5.81)
whenever $2\beta\leq\theta$. Also, for any $k\in\mathbb{N}$, the terms $|\partial_{t}^{k}\widehat{E}_{0}(t,\xi)|$ and $|\partial_{t}^{k}\widehat{E}_{1}(t,\xi)|$ can be estimated as
$|\partial_{t}^{k}\widehat{E}_{1}(t,\xi)|$ $\lesssim\lambda_{\xi}^{(k-1)\theta}e^{-t},$ (5.82)
and
$|\partial_{t}^{k}\widehat{E}_{0}(t,\xi)|$ $\lesssim\lambda_{\xi}^{k\theta}e^{-t}.$ (5.83)
Estimate for $\|u(t)\|_{\mathcal{H}}$: Recalling the Plancherel formula along with the estimates (5.20), (5.71), (5.75), and (5.78), we have
$\|u(t)\|_{\mathcal{H}}^{2}\lesssim\sum_{\xi\in\mathcal{R}_{1}}\left(|\widehat{u}_{0}(\xi)|^{2}+t^{2}|\widehat{u}_{1}(\xi)|^{2}\right)+\sum_{\xi\in\mathcal{R}_{2}}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)\\ +\sum_{\xi\in\mathcal{R}_{3}}(1+t)^{2}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)+\sum_{\xi\in\mathcal{R}_{4}}e^{-2t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)\lesssim\|u_{0}\|^{2}_{\mathcal{H}}+t^{2}\|u_{1}\|^{2}_{\mathcal{H}}.$ (5.84)
$+\sum_{\xi\in\mathcal{R}_{4}}e^{-2t}\left(\lambda_{\xi}^{2\beta}|\widehat{u}_{0}(\xi)|^{2}+\lambda_{\xi}^{2\beta-\theta}|\widehat{u}_{1}(\xi)|^{2}\right)\lesssim e^{-\delta t}\begin{cases}\left(\|u_{0}\|^{2}_{\mathcal{H}^{2\beta}_{\mathcal{L}}}+\|u_{1}\|^{2}_{\mathcal{H}^{2\beta-\theta}_{\mathcal{L}}}\right)&\text{if }2\beta>\theta,\\ \left(\|u_{0}\|^{2}_{\mathcal{H}^{2\beta}_{\mathcal{L}}}+\|u_{1}\|^{2}_{\mathcal{H}}\right)&\text{if }2\beta\leq\theta.\end{cases}$ (5.85)
Estimate for $\|\partial_{t}^{k}u(t)\|_{\mathcal{H}}$: Combining (5.22), (5.74), (5.77), (5.82), and (5.83), we get
$\|\partial_{t}u(t)\|_{\mathcal{H}}^{2}$ $\lesssim\sum_{\xi\in\mathcal{R}_{1}}|\widehat{u}_{1}(\xi)|^{2}+\sum_{\xi\in\mathcal{R}_{2}}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)+\sum_{\xi\in\mathcal{R}_{3}}(1+t)^{2}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)$
$+\sum_{\xi\in\mathcal{R}_{4}}e^{-2t}\left(\lambda_{\xi}^{2\theta}|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)\lesssim e^{-\delta t}\|u_{0}\|^{2}_{\mathcal{H}^{2\theta}_{\mathcal{L}}}+\|u_{1}\|^{2}_{\mathcal{H}}.$ (5.86)
Furthermore, for $k\geq 2$, we have
$\|\partial_{t}^{k}u(t)\|_{\mathcal{H}}^{2}$ $\lesssim\sum_{\xi\in\mathcal{R}_{2}}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)+\sum_{\xi\in\mathcal{R}_{3}}(1+t)^{2}e^{-\delta t}\left(|\widehat{u}_{0}(\xi)|^{2}+|\widehat{u}_{1}(\xi)|^{2}\right)$
$+\sum_{\xi\in\mathcal{R}_{4}}e^{-2t}\left(\lambda_{\xi}^{2k\theta}|\widehat{u}_{0}(\xi)|^{2}+\lambda_{\xi}^{(2k-1)\theta}|\widehat{u}_{1}(\xi)|^{2}\right)\lesssim e^{-\delta t}\left(\|u_{0}\|^{2}_{\mathcal{H}^{2k\theta}_{\mathcal{L}}}+\|u_{1}\|^{2}_{\mathcal{H}^{2(k-1)\theta}_{\mathcal{L}}}\right).$ (5.87)
@@ 6. Application
Having established the decay estimates for both the undamped and damped cases of the semilinear Cauchy problem (1.3), we are now able to employ these results to the question of global/local (in-time) well-posedness. This application emphasizes the decay framework’s efficiency in managing nonlinear effects and guaranteeing the long-term existence and uniqueness of solutions under appropriate assumptions regarding the initial data and the nonlinearity.
We can recognize our Hilbert space $\mathcal{H}$ with $L^{2}(\Omega)$ for some measure space $\Omega$, thereby enabling us to utilize the entire scale of $L^{p}(\Omega)$ spaces. An exemplary case of interest is $\mathcal{H}=L^{2}(\mathbb{R}^{n})$ or $\mathcal{H}=L^{2}(\mathcal{M})$ for a compact manifold $\mathcal{M}$. In this notation, we denote $\|\cdot\|_{\mathcal{H}}=\|\cdot\|_{2}$ and $\|\cdot\|_{p}=\|\cdot\|_{L^{p}(\Omega)}$.
Definition 6.1 (Gagliardo–Nirenberg index).
For the operator $\mathcal{L}$, we say that $p\geq 1$ is Gagliardo Nirenberg admissible if the following Gagliardo–Nirenberg type inequality
$\|u\|_{2p}\lesssim\|\mathcal{L}^{\sigma/2}u\|^{\alpha}_{2}\|u\|_{2}^{1-\alpha}$ (6.1)
holds for some $\alpha=\alpha(p)\in[0,1]$.
Remark 6.2.
Here are some examples of Gagliardo–Nirenberg admissible indices:
  • For harmonic sscillator $\mathcal{L}=-\Delta+|x|^{2}$ on $L^{2}(\mathbb{R}^{n})$ or for Laplacian on $L^{2}(\mathbb{S}^{n})$: The Gagliardo–Nirenberg admissible indices are given by
    $\begin{cases}n=1\text{ and }n=2&:~~1\leq p<\infty,\\ n\geq 3&:~~1\leq p\leq\frac{n}{n-2\sigma}.\end{cases}$ (6.2)
  • Laplacian on compact Riemannian manifolds: The Gagliardo–Nirenberg admissible indices are given by
    $\begin{cases}n=2&:~~1\leq p<\infty,\\ n\geq 3&:~~1\leq p\leq\frac{n}{n-2\sigma}.\end{cases}$ (6.3)
  • Twisted Laplacian (Landau Hamiltonian) on $\mathbb{C}^{n}$ : The Gagliardo–Nirenberg admissible indices are given by
    $\begin{cases}n=1&:~~1\leq p<\infty,\\ n\geq 2&:~~1\leq p\leq\frac{n}{n-\sigma}.\end{cases}$ (6.4)
For more details regarding the examples in the above remark, one can refer to [mic-wave2, Section 3] and the references therein.
@@
Let $\sigma>0$, $\theta\in[0,\sigma]$ and $\mathcal{L}$ be strictly positive operator. Let $p\geq 1$ be Gagliardo-Nirenberg admissible according to Definition 6.1 and $f$ satisfies the following relation:
$|f(u)-f(v)|\lesssim(|u|^{p-1}+|v|^{p-1})|u-v|\quad\text{and }\quad f(0)=0.$ (6.5)
Also assume that
$(u_{0},u_{1})\in\mathcal{A}^{\sigma,\theta}:=\begin{cases}\mathcal{H}_{\mathcal{L}}^{\sigma}\times\mathcal{H}\quad&\mathrm{for}~~2\theta\in[0,\sigma]\\ \mathcal{H}_{\mathcal{L}}^{2\theta}\times\mathcal{H}\quad&\mathrm{for}~~2\theta\in(\sigma,2\sigma].\end{cases}$
@@
$(u_{0},u_{1})\in\mathcal{A}^{\sigma,\theta}~~\text{ satisfying }~~\|(u_{0},u_{1})\|_{\mathcal{A}^{\sigma,\theta}}<\epsilon,$
there is a unique global solution $u\in C(\mathbb{R}_{+};\mathcal{H}_{\mathcal{L}}^{\sigma})\cap C^{1}(\mathbb{R}_{+};\mathcal{H})$ to the Cauchy problem (1.3). Moreover, $u$ also satisfies the estimates given in Theorem 2.5 for $(k,\beta)\in\{(0,0),(0,\sigma/2),(1,0)\}$ and for all $t>0$.
As mentioned in the prior discussion, we denote $\|\cdot\|_{\mathcal{H}}=\|\cdot\|_{2}$.
@@
represents the solution of the homogeneous Cauchy problem (2.1), where $_{\mathcal{L}}$ denotes $\mathcal{L}$-convolution. The solution to the semilinear Cauchy problem (1.3), using Duhamel’s principle, can be expressed as
@@
Now, we introduce the solution space $X:=C(\mathbb{R}_{+};\mathcal{H}_{\mathcal{L}}^{\sigma})\cap C^{1}(\mathbb{R}_{+};\mathcal{H})$ associated with the norm
$\left\|u\right\|_{X}:=\sup\limits_{t\in[0,\infty)}\left\{e^{\delta t}\left(\left\|u(t)\right\|_{2}+\left\|\mathcal{L}^{\sigma/2}u(t)\right\|_{2}+\left\|\partial_{t}u(t)\right\|_{2}\right)\right\}.$
Let $\mathcal{N}:X\to X$ be the operator defined as follows: